candice is four years older than rosa, and the sum of their ages is 22. How old is rosa?
Found 2 solutions by stanbon, lynnlo:
Answer by stanbon(75887)

candice is four years older than rosa, and the sum of their ages is 22. How old is rosa?
----
Equations:
c = r + 4
c + r = 22
----
Substitute for "c" and solve for "r":
r+4 + r = 22
2r = 18
r = 9 (Rosa's age now)
===========================
Cheers,
Stan H.

Answer by lynnlo(4176)

candice is 13
rosa is 9
13+9=22